# Neural Spelling: A Spell-Based BCI System for Language Neural Decoding

### v1

@@ Abstract
Brain-computer interfaces (BCIs) present a promising avenue by translating neural activity directly into text, eliminating the need for physical actions. However, existing non-invasive BCI systems have not successfully covered the entire alphabet, limiting their practicality. In this paper, we propose a novel non-invasive EEG-based BCI system with Curriculum-based Neural Spelling Framework, which recognizes all 26 alphabet letters by decoding neural signals associated with handwriting first, and then apply a Generative AI (GenAI) to enhance spell-based neural language decoding tasks. Our approach combines the ease of handwriting with the accessibility of EEG technology, utilizing advanced neural decoding algorithms and pre-trained large language models (LLMs) to translate EEG patterns into text with high accuracy. This system show how GenAI can improve the performance of typical spelling-based neural language decoding task, and addresses the limitations of previous methods, offering a scalable and user-friendly solution for individuals with communication impairments, thereby enhancing inclusive communication options.
Index Terms:
Fuzzy Logic, Transformers, fNIRS, Social Neuroscience
@@ I Introduction
Brain-computer interfaces (BCIs) have emerged as a pivotal area of research within human-computer interaction (HCI), distinguished by their capacity to seamlessly integrate neural signals with external systems. Pioneering studies such as those by Guo et al.[1], Chen et al.[2], Cao et al.[3], and Lin et al.[4] underscore BCIs’ role in advancing neuroscience and technology. These interfaces create direct communication pathways that are especially beneficial for individuals with limited speech or motor functions. Language decoding represents a critical domain within BCI research, aimed at deciphering the neural correlates of speech and language processing. This capability not only enhances interaction paradigms but also opens new avenues for communication, offering significant improvements in quality of life for those with severe communicative impairments.
Figure 1: Demonstrations of two typical language decoding frameworks. (A) Direct speech synthesis approach. (B) Spelling-based approach using phonemes. (C) Direct text synthesis approach. (D) Spelling-based approach using letters.
Initially, BCI research focused on visual-based decoding approaches, such as steady-state visually evoked potentials (SSVEP)[1], which, despite their reliability, demand high cognitive effort and are unsuitable for prolonged use[2]. These methods often fail to align with natural human language patterns, posing significant usability challenges. The advent of invasive neural decoding technologies marked a significant advancement, allowing for direct interpretation of brain signals via electrocorticography (ECoG) or stereo-EEG (sEEG)[5, 6]. These methods have demonstrated substantial improvements in user performance, significantly enhancing communicative capacities for patients with speech impairments. However, the invasive nature, high cost, and ethical concerns limit their general applicability[7]. In contrast, non-invasive BCIs, predominantly utilizing electroencephalography (EEG), offer a more accessible alternative. These systems are less obtrusive and more cost-effective, broadening potential user demographics [8]. Despite the challenges of signal noise and the extensive training required for users, recent studies have demonstrated EEG’s potential in effective language decoding [9, 10].
@@
Figure 2: (A): The Experiment Design. (B): The trajectory finishing with time. (C): The PSD feature calculated from Raw EEG signal. (D): The structures of the Trajectory Resnet18 Encoder and the Converlutional EEG Encoder. (E): The Letter Probability Distribution in the outputs from the classifier layers. (F): The structure of the sentence generator.
@@
[…]
@@
This paper contributes to the evolving landscape of brain-computer interfaces (BCIs) by proposing a hybrid approach that merges the accessibility of non-invasive EEG with the advanced linguistic capabilities of GenAI. We introduce a novel curriculum-based neural spelling framework (CNS) for neural language decoding. This framework employs a convolutional neural network (CNN)-based encoder to initially learn individual-specific letter transition patterns, followed by a curriculum-driven LLM to synthesize sentence texts. The results demonstrate significant enhancements in the performance of language decoding tasks. By integrating these advanced technologies, we aim to facilitate more natural, efficient, and inclusive communication methods for individuals with varying physical abilities, heralding a new era in assistive communication technologies. This introduction outlines the transformation of BCI technology from basic visual decoding to sophisticated language synthesis, setting the stage for a detailed discussion on how current technologies can be integrated and enhanced to better meet the communicative needs of a diverse user population.
@@ IV-A1 Stage 1: Neural Letter Classifier
The Convolutional EEG Encoder is designed as a two-layer Convolutional Neural Network (CNN), which is adept at processing EEG data for the purpose of embedding generation, as shown in Fig. 2D. This encoder operates in conjunction with CL techniques to enhance the discriminative power of the embeddings. The trajectory data is processed by a pre-trained ResNet18 [48] model, which serves as a trajectory encoder. The core of our learning strategy is based on minimizing the contrastive learning loss ($loss_{CL}$), which is calculated as follows:
@@
$loss_{CL}=1-\cos(\theta_{EEG},\theta_{Traj})^{2}$ (3)
@@
The letter classification component is an integral part of the CNS Framework. It includes a fully connected linear layer that maps its input to a 26-node output layer, where each node represents a letter of the English alphabet. After this. This arrangement enables the system to predict letters based on the processed EEG data. A softmax function is applied to the output layer to obtain a probability distribution over the alphabet, as shown in Fig. 2E, which is critical for the system’s accuracy in spelling prediction. The classifier utilizes the cross-entropy loss to measure the discrepancy between the predicted probabilities and the actual distribution of the target letters. The crossentrypy loss ($loss_{CE}$) function is defined as follows:
@@
where $M$ is the number of classes (letters), $y_{o,c}$ is a binary indicator (0 or 1) if class label $c$ is the correct classification for observation $o$, and $p_{o,c}$ is the predicted probability of observation $o$ being of class $c$. This loss function effectively guides the learning process by penalizing deviations from the true label distribution, thus enhancing the system’s ability to generate accurate and reliable spelling predictions from EEG signals.
In training phase, the finial loss ($loss_{total}$) is calculated by:
$loss_{total}=0.35\times loss_{CE}+0.65\times loss_{CL}$ (5)
@@ IV-A2 Stage 2: Curriculum Learning for Language Model Integration in Fluent Sentence Generation
Ground Truth: This civilization, known as the ’Ayar,’ possesses an advanced understanding of nature and the power to manipulate the very fabric of reality.
@@
where $N_{\omega}$ denotes the count of samples for letter $\omega$. From this, the letter $\omega$ is then probabilistically sampled based on its presence within the Top-$K$ distribution. For instance, consider the ground truth (GT) sentence, ”As he chases down rogue time travelers…”. If the aggregated probability for ’A’ at the start, $C^{*}_{A}$, is 0.2 and for ’R’, $C^{*}_{R}$, is 0.6, the classifier might predict ”Rs” in place of ”As”.
@@
In this phase, we incorporate the noisy predictions from the Neural Letter Classifier (NLC) into the LLM’s training regimen. Specifically, we replace selected characters in the training texts with their corresponding noisy variants from the Top-$K$ output distributions. The LLM is then fine-tuned to de-noise this altered text. This structured noise adaptation process is designed to enhance the model’s robustness, enabling it to effectively manage and correct noisy inputs during inference.
@@
We employ bi-directional attention during training for this seq2seq generation task, as shown in Fig. 2(F). The fine-tuning objective combines the denoising training objective with the causal language modeling objective [45], facilitating the model’s use of previously generated tokens to guide the prediction of subsequent tokens, thereby reducing the influence of noisy letter-level inputs.
@@ V-A Participants
We recruited thirty-two right-handed, healthy individuals (P1–P32), all native English speakers from Australia, comprising 16 males and 16 females with an average age of $M_{\text{age}}=24.94\pm 0.29$ years. Each participant had normal or corrected-to-normal vision and reported no history of mental health disorders. 4 subjects are removed from the dataset due to their misunderstanding of the experiment design. Ethical approval was granted by the University of Technology Sydney’s Ethics Committee (Approval No. UTS HREC REF: ETH23-8036). Informed consent was obtained in writing from each participant prior to the experiment.
Figure 3: Model Comparisons: Top 1, Top 3, and Top 5 Accuracy of Different Models. Orange color comparisons shows the difference between w/CL and w/o CL conditions. Asterisks (***) indicate significant results ($\textit{p}~{}<~{}0.001$).
@@ V-B Instrumentation
EEG signals were recorded using a 64-channel Neuroscan amplifier (Curry 9), with electrode placements following the 10–20 international system [49]. A ground electrode was referenced to maintain signal consistency.
@@ V-C Experimental Design
Figure 4: Top 1 Accuracy of CNN w/ CL Model Across Different ROIs and Frequency Bands.
@@ V-D Data Processing
Figure 5: (A): The KDE distributions among only neural letter classifier (NLC), only LLMs (baseline), and LLMs after NLC. (B): The linear regression between the CER and WER scores of NLC ($S_{NLC}$) and LLMs ($S_{LLMs}$).
@@
EEG data preprocessing was performed using MNE (version 1.6.0) [50] in Python 3.10.13. The machine learning pipeline began with a bandpass filter between 1 and 70 Hz to reduce general noise. This was followed by notch filtering at 50, 100, and 150 Hz to mitigate interference from power lines. Next, epochs were created, defining time intervals from -1 s to 3 s around each event of the first pen-down in each trial. Independent Component Analysis (ICA) was applied with the auto-reject method [51] to remove artifacts from the data. Finally, the EEG data was re-referenced by averaging and baseline correction was applied to ensure consistency across channels.
@@
Power Spectral Density (PSD) was extracted as the primary feature for model prediction. To compute PSD, an Fast Fourier Transform (FFT) was used, capturing frequencies between 1 Hz and 70 Hz, as shown in Fig. 2C. The PSD for a given signal $x(t)$ was calculated using the formula:

### v2

# arXiv:2501.17489v2[cs.HC] 02 Mar 2026

@@ Abstract
Objective: Brain–computer interfaces (BCIs) can support communication by translating neural activity into text, yet existing non-invasive systems rarely cover the full alphabet, limiting their practical use. Methods: We propose a novel non-invasive EEG-based BCI framework, Curriculum-based Neural Spelling (CNS), that decodes all 26 English letters by first learning neural patterns associated with handwriting trajectories. A Generative AI (GenAI) module based on large language models (LLMs) is then integrated to refine spell-based neural decoding and correct semantic errors. Results: The proposed system achieves robust letter-level decoding and accurate sentence reconstruction across users, outperforming conventional EEGNet and hybrid CNN–RNN baselines. GenAI correction further reduces word error rates and enhances decoding fluency. Conclusion: Combining EEG-based neural spelling with generative language modeling enables full-alphabet decoding and significantly improves linguistic accuracy in non-invasive BCIs. Significance: This work demonstrates how integrating GenAI with neural decoding can bridge the gap between noisy signal-level predictions and coherent language-level outputs, establishing a system-level framework for full-alphabet neural spelling and adaptive language-level correction under non-invasive EEG constraints.
@@ I Introduction
Brain-computer interfaces (BCIs) have emerged as a pivotal area of research within human-computer interaction (HCI), distinguished by their capacity to seamlessly integrate neural signals with external systems. Pioneering studies such as those by Guo et al.[26], Chen et al.[9], Cao et al.[7], Nguyen et al.[53], and Lin et al.[42] underscore BCIs' role in advancing neuroscience and technology. These interfaces create direct communication pathways that are especially beneficial for individuals with limited speech or motor functions. Language decoding represents a critical domain within BCI research, aimed at deciphering the neural correlates of speech and language processing. This capability not only enhances interaction paradigms but also opens new avenues for communication, offering significant improvements in quality of life for those with severe communicative impairments.
Figure 1: Demonstrations of four typical language decoding frameworks. (A) Direct speech synthesis approach. (B) Spelling-based approach using phonemes. (C) Direct text synthesis approach. (D) Spelling-based approach using letters.
Initially, BCI research focused on visual-based decoding approaches, such as steady-state visually evoked potentials (SSVEP)[26], which, despite their reliability, demand high cognitive effort and are unsuitable for prolonged use[9]. These methods often fail to align with natural human language patterns, posing significant usability challenges. The advent of invasive neural decoding technologies marked a significant advancement, allowing for direct interpretation of brain signals via electrocorticography (ECoG) or stereo-EEG (sEEG)[3, 2]. These methods have demonstrated substantial improvements in user performance, significantly enhancing communicative capacities for patients with speech impairments. However, the invasive nature, high cost, and ethical concerns limit their general applicability[76]. In contrast, non-invasive BCIs, predominantly utilizing electroencephalography (EEG), offer a more accessible alternative. These systems are less obtrusive and more cost-effective, broadening potential user demographics [66, 29]. Despite the challenges of signal noise and the extensive training required for users, recent studies have demonstrated EEG's potential in effective language decoding [13, 73].
@@
Figure 2: (A): The experimental design. (B): The trajectory completion over time. (C): The power spectral density (PSD) features computed from raw EEG signals. (D): The structures of the Trajectory ResNet-18 encoder and the Convolutional EEG encoder. (E): The letter probability distribution at the classifier output. (F): The structure of the sentence generator.
@@
The spell-based methods initially utilize N-gram models to calculate the probability of sequential letter occurrences. For instance, the ECoG study aimed at spelling imaged handwritten letters [75], and the sEEG study focused on spelling the Pinyin representations of Chinese pronunciations [21]. Subsequently, GenAI technologies have been employed to enhance performance, as demonstrated in studies like [76] and [50]. These vocal-based methods underscore the potential of GenAI for neural language decoding. However, these approaches have limitations. For example, vocal-based methods require accurate pronunciation by subjects. Given the diversity in pronunciation—such as the thousands of syllables and 44 phonemes in English, including closely related sounds like // and /ae/, along with variations between short and long vowels—it becomes challenging to accurately generate complete sentences. Simultaneously, Mandarin Chinese combines 21 consonant phonemes, 7 vowel phonemes, and 4 tones to create over 400 distinct syllables [21]. This vast array of phonetic elements significantly complicates speech synthesis in both languages. The complexity is further exacerbated in non-invasive speech synthesis applications due to the signal-to-noise ratio (SNR) and the limitations of the training datasets available. These factors together pose substantial challenges in developing robust and accurate speech synthesis systems. Therefore, spelling letters, as opposed to using vocal-based features, presents a more effective option for language decoding. The classification of just 26 letters not only simplifies the process compared to pronunciation-based methods but also enhances performance and applicability, even facilitating cross-linguistic utility, such as seamlessly integrating English and Pinyin. Unlike prior EEG-based handwriting paradigms that typically focus on limited symbol sets or isolated character recognition, the present study systematically investigates full-alphabet (26-letter) decoding and further integrates curriculum-adapted language modeling to enable sentence-level reconstruction. To optimize letter input, innovative scenarios have been designed, such as 'air-writing' [67] and 'paper-writing' [55].
Moreover, cVEP-based systems have demonstrated efficient and reliable performance under non-invasive EEG settings [22], and online P300 spellers have been established as practical communication systems since the seminal work of Farwell and Donchin [20]. More recently, complete sentence-level online spelling pipelines with calibration procedures and large language model integration have also been reported [30]. These approaches represent well-established non-invasive BCI paradigms.
However, a common characteristic of these systems is that they are predominantly stimulus-driven or reactive in nature, relying on externally evoked neural responses (e.g., visual flickers or oddball paradigms) and predefined symbol sets. While such paradigms achieve high information transfer rates and robust online performance, they are primarily designed to optimize communication efficiency, but this comes at the cost of increased visual fatigue.
In contrast, handwriting-based BCIs aim to decode active, self-initiated motor-related neural activity associated with voluntary letter production, enabling more naturalistic and flexible letter generation without reliance on continuous external stimulation. Despite extensive research on non-invasive spellers, whether an active, handwriting-based BCI can reliably decode the full alphabet (26 letters) from scalp EEG, and thereby establish a normative neural reference for neural handwriting under realistic non-invasive constraints, remains an underexplored problem. The present work addresses this gap by focusing on neural handwriting as an active input paradigm, rather than on stimulus-driven letter selection or communication throughput alone.
@@
This paper contributes to the evolving landscape of brain-computer interfaces (BCIs) by proposing a hybrid approach that merges the accessibility of non-invasive EEG with the advanced linguistic capabilities of GenAI. We introduce a novel curriculum-based neural spelling framework (CNS) for neural language decoding. This framework employs a convolutional neural network (CNN)-based encoder to initially learn individual-specific letter transition patterns, followed by a curriculum-driven LLM to synthesize sentence texts. The results demonstrate significant enhancements in the performance of language decoding tasks. By integrating these advanced technologies, we establish a normative neural baseline for non-invasive neural handwriting under realistic EEG constraints. We aim to facilitate a system-level investigation of full-alphabet neural spelling, demonstrating how low-SNR EEG decoding can be coupled with curriculum-adapted generative language models to enable language-level error correction, heralding a new era in assistive communication technologies. This introduction outlines the transformation of BCI technology from basic visual decoding to sophisticated language synthesis, setting the stage for a detailed discussion on how current technologies can be integrated and enhanced to better meet the communicative needs of a diverse user population.
@@ IV-A1 Stage 1: Neural Letter Classifier
The Convolutional EEG Encoder is designed as a two-layer Convolutional Neural Network (CNN), which is adept at processing EEG data for the purpose of embedding generation, as shown in Fig. 2D. This encoder operates in conjunction with contrastive learning (ConL) techniques to enhance the discriminative power of the embeddings. The trajectory data is processed by a pre-trained ResNet18 [28] model, which serves as a trajectory encoder. The core of our learning strategy is based on minimizing the contrastive learning loss ($loss_{{\color[rgb]{0,0,0}ConL}}$), which is calculated as follows:
@@
$loss_{{\color[rgb]{0,0,0}ConL}}=1-\cos(\theta_{EEG},\theta_{Traj})^{2}$ (3)
@@
The letter classification component is an integral part of the CNS Framework. It includes a fully connected linear layer that maps its input to a 26-node output layer, where each node represents a letter of the English alphabet. After this. This arrangement enables the system to predict letters based on the processed EEG data. A softmax function is applied to the output layer to obtain a probability distribution over the alphabet, as shown in Fig. 2E, which is critical for the system's accuracy in spelling prediction. The classifier utilizes the cross-entropy loss to measure the discrepancy between the predicted probabilities and the actual distribution of the target letters. The crossentrypy loss ($loss_{CE}$) function is defined as follows:
@@
where $M$ is the number of classes (letters), $y_{o,c}$ is a binary indicator (0 or 1) if class label $c$ is the correct classification for observation $o$, and $p_{o,c}$ is the predicted probability of observation $o$ being of class $c$. This loss function effectively guides the learning process by penalizing deviations from the true label distribution, thus enhancing the system's ability to generate accurate and reliable spelling predictions from EEG signals.
In the training phase, the final loss ($loss_{total}$) is computed as a weighted combination of the cross-entropy loss ($loss_{CE}$) and the contrastive loss ($loss_{{\color[rgb]{0,0,0}ConL}}$):
$loss_{total}=0.35\times loss_{CE}+0.65\times loss_{{\color[rgb]{0,0,0}ConL}}.$ (5)
The weighting coefficients are selected empirically to balance classification accuracy and representation regularization.
@@ IV-A2 Stage 2: Curriculum Learning for Language Model Integration in Fluent Sentence Generation
Ground Truth: This civilization, known as the 'Ayar,' possesses an advanced understanding of nature and the power to manipulate the very fabric of reality.
@@
where $N_{\omega}$ denotes the count of samples for letter $\omega$. From this, the letter $\omega$ is then probabilistically sampled based on its presence within the Top-$K$ distribution. For instance, consider the ground truth (GT) sentence, "As he chases down rogue time travelers…". If the aggregated probability for 'A' at the start, $C^{*}_{A}$, is 0.2 and for 'R', $C^{*}_{R}$, is 0.6, the classifier might predict "Rs" in place of "As".
@@
Noise adaptation refers to the process of fine-tuning the language model on EEG decoder outputs that contain systematic letter-level errors, enabling the model to recognize and correct typical noise patterns arising from neural decoding during inference. In this phase, we incorporate the noisy predictions from the Neural Letter Classifier (NLC) into the LLM's training regimen. Specifically, we replace selected characters in the training texts with their corresponding noisy variants from the Top-$K$ output distributions. The LLM is then fine-tuned to de-noise this altered text. This structured noise adaptation process is designed to enhance the model's robustness, enabling it to effectively manage and correct noisy inputs during inference.
@@
We employ bi-directional attention during training for this seq2seq generation task, as shown in Fig. 2(F). The fine-tuning objective combines the denoising training objective with the causal language modeling objective [39], facilitating the model's use of previously generated tokens to guide the prediction of subsequent tokens, thereby reducing the influence of noisy letter-level inputs.
@@
The sensitivity analyses of the proposed curriculum learning (CL) strategy with respect to two key factors, noise injection strength and the number of curriculum stages, as well as the influence of different curriculum schedule shapes on model performance, are presented in Supplemental Sections I.A and I.B.
@@ V-A Participants
We recruited thirty-two right-handed, healthy individuals (P1–P32), all native English speakers from Australia, comprising 16 males and 16 females with an average age of $M_{\text{age}}=24.94\pm 0.29$ years. Each participant had normal or corrected-to-normal vision and reported no history of mental health disorders. Four subjects were removed from the dataset due to misunderstanding of the experimental instructions. An a priori power analysis (G*Power v3.1.9.7) indicated that approximately 30 participants are sufficient to detect medium within-subject effects (Cohen’s $d_{z}=0.5$) at a significance level of $\alpha=0.05$. The final sample size therefore meets standard statistical power requirements for within-subject EEG analyses. Ethical approval was granted by the University of Technology Sydney's Ethics Committee (Approval No. UTS HREC REF: ETH23-8036). Informed consent was obtained in writing from each participant prior to the experiment.
@@ V-B Instrumentation
EEG signals were recorded using a 64-channel Neuroscan amplifier (Curry 9; Australia), with electrode placements following the 10–20 international system [34]. A ground electrode was referenced to maintain signal consistency.
@@ V-C Experimental Design
Figure 3: Top 1 Accuracy of CNN w/ ConL Model Across Different ROIs and Frequency Bands.
@@ V-D Data Processing
Figure 4: (A): The Kernel Density Estimate (KDE) distributions among only neural letter classifier (NLC), only LLMs (Large Language Models, the baseline), and LLMs after NLC. (B): The linear regression between the CER and WER scores of NLC ($S_{NLC}$) and LLMs ($S_{LLMs}$).
@@
EEG data preprocessing was performed using MNE (version 1.6.0) [25] in Python 3.10.13. The pipeline began with a bandpass filter between 1 and 70 Hz to attenuate slow drifts and high-frequency noise, followed by notch filtering at 50, 100, and 150 Hz to suppress line noise and its harmonics. EEG signals were then epoched from 1-1 s to $3$ s relative to the first pen-down event in each trial, capturing preparatory, execution, and post-execution neural activity. Artifact handling was performed in two stages. First, automated channel-level artifact suppression was applied using the Autoreject framework [33] to detect and interpolate noisy channels within each epoch; no trials were discarded at this stage. This step improves data quality for subsequent analysis while preserving the full trial structure. Second, Independent Component Analysis (ICA) was applied to the epoched data after channel interpolation. ICA decomposition was performed using the FastICA algorithm implemented in MNE. Independent components were inspected following established EEG practice based on their temporal dynamics, spatial topographies, and spectral characteristics. Only clearly identifiable ocular artifacts (e.g., blink-related components and other clearly non-neural artifacts) were removed using a conservative manual procedure, with fewer than three components removed per subject. Finally, the EEG data were re-referenced to the common average, and baseline correction was applied using the pre-event interval to ensure consistency across channels.
TABLE II: Paired comparison between the proposed CNN-based model and representative baselines, with and without contrastive learning (ConL). Accuracy is reported as mean $\pm$ standard deviation (%). Statistical significance is assessed using paired t-tests across subject-sessions, with $p<0.05$, ∗∗ $p<0.01$, and ∗∗∗ $p<0.001$, comparing each baseline against the proposed CNN model.
Model Setting Top-1 Top-3 Top-5
CNN (Proposed) w/o ConL $27.23\pm 12.10$ $51.34\pm 14.78$ $65.29\pm 12.08$
w/ ConL $33.10\pm 11.54$ $57.77\pm 12.89$ $71.46\pm 11.11$
LSTM w/o ConL $18.22\pm 3.46^{***}$ $31.21\pm 6.71^{***}$ $41.66\pm 6.42^{***}$
w/ ConL $25.74\pm 3.51^{**}$ $37.85\pm 4.36^{***}$ $49.00\pm 6.07^{***}$
Transformer w/o ConL $15.54\pm 4.40^{***}$ $27.56\pm 4.14^{***}$ $36.44\pm 4.02^{***}$
[…]
EEG-Deformer w/o ConL $17.94\pm 4.46^{***}$ $30.02\pm 4.12^{***}$ $38.82\pm 4.04^{***}$
w/ ConL $28.99\pm 3.14^{*}$ $38.22\pm 3.00^{***}$ $48.92\pm 3.50^{***}$
Conformer w/o ConL $14.59\pm 4.57^{***}$ $26.54\pm 4.19^{***}$ $35.28\pm 3.73^{***}$
w/ ConL $25.56\pm 3.29^{**}$ $34.37\pm 3.58^{***}$ $45.54\pm 3.22^{***}$
EEGFormer w/o ConL $15.21\pm 4.97^{***}$ $26.87\pm 4.67^{***}$ $35.74\pm 3.74^{***}$
w/ ConL $26.07\pm 3.38^{**}$ $35.16\pm 3.19^{***}$ $46.46\pm 3.86^{***}$
TCNet w/o ConL $16.34\pm 4.63^{***}$ $28.64\pm 4.28^{***}$ $37.46\pm 3.96^{***}$
w/ ConL $27.53\pm 3.17^{**}$ $36.82\pm 3.30^{***}$ $47.17\pm 3.60^{***}$
@@
Power Spectral Density (PSD) was extracted as the primary feature for model prediction. To compute PSD, an Fast Fourier Transform (FFT) was used capturing frequencies between 1 Hz and 70 Hz, as shown in Fig. 2C. The PSD for a given signal $x(t)$ was calculated using the formula: